huilan is 14 years older than thomas. the sum of their ages is 84. what is thomas's age
Found 2 solutions by ikleyn, Alan3354:
Answer by ikleyn(52787)

.
Huilan + Thomas = 84


(T+14) + T      = 84


2T + 14 = 84


2T = 84 - 14 = 70


T = 70/2 = 35.


Answer.  Thomas is 35 years old.  Huilan is 35 + 14 = 49 years old.

Solved.

Do you understand everything as it is presented in my solution ?



Answer by Alan3354(69443)

h - t = 14
h + t = 84
------------- Add